OpenStudy (e.mccormick):

The MathJax-\(\LaTeX\) Chronicles A introduction to MathJax-\(\LaTeX\) that will include some video segments.

OpenStudy (e.mccormick):

\(\TeX\) is a cross platform document typesetting system initially made in 1978. In 1984, the \(\LaTeX\) project extended this with a more high level language to try and make it more accessible. In 2004 the jsMath project started using these older typesetting languages as a foundation for showing math on the Internet. Today, MathJax, the successor to jsMath, is used on OpenStudy to convert \(\LaTeX\) codes into properly formatted math, art, and much more. MathJax does not support all of \(\LaTeX\), but it does let people use a very wide range of it when they do their math, art, etc. The first thing you need to realize about the \(\TeX\) macros you can type here is that most of them are named after what they do. This makes remembering them a bit easier because they have a relationship to what you are doing. For example, `\frac{}{}` makes a fraction and `\sqrt{}` makes a square root. The second thing you need to realize is that not all \(\TeX\) is needed or even works here. If you try to get information from a \(\LaTeX\) help site or video, it might show you a lot more than what you need! Now, lets look at doing a little \(\LaTeX\) on OpenStudy. First, I want to show three ways to do the Quadratic Formula: the OS Equation Editor, the Code Cogs Equation Editor, and by hand. \(x=\dfrac{-b\pm\sqrt{b^2-4ac}}{2a}\) http://youtu.be/6vML7NyKscE

OpenStudy (e.mccormick):

As I did that, you saw I used `\[ \]` around some of it and `\( \)` around other parts. These are the block mode and inline (text) mode for the MathJax engine. Typically I use the inline mode and ignore blocks. Here is a video so you can see the preview of what is happening as I type the below: http://youtu.be/qz7FfrS-6lo \[f'(a)=\lim_{h\rightarrow 0} \dfrac{f(a+h)-f(a)}{h}\] \[f\,\prime(a)=\lim_{h\rightarrow 0} \dfrac{f(a+h)-f(a)}{h}\] \(f\,\prime(a)=\lim_{h\rightarrow 0} \dfrac{f(a+h)-f(a)}{h}\) \(f\,\prime(a)=\lim\limits_{h\rightarrow 0} \dfrac{f(a+h)-f(a)}{h}\) \[\sum_{i=0}^7i^2\] What about inline mode? \(\sum\limits_{i=0}^7i^2\) Why is that nice at times? I am going to pad this paragraph some with a bit of useless text to make it wrap a couple lines. That way you can see how the two different inline parts, the one with \limits and the one without it are going to end up with different line spacings. \(\sum_{i=0}^7i^2\) This is a longer paragraph. This is a longer paragraph. This is a longer paragraph. And that should be long enough.

OpenStudy (e.mccormick):

The next topic is the \(\textbf{artistic use}\) of \(\LaTeX\). While this all started as a way to format math, it has allowed a lot more formatting to be done. Some of this is very useful or artistic. An example is the use of the words in bold above. Another would be to add color: \(\color{red}{\text{red color}}\) or a bounding box, \(\bbox[orange, 1px,border:2px solid black]{\color{blue}{\mathfrak{BOO!}}}\). There is a lot more to say on this topic, but I will leave that up to the assorted tutorials. I just want to show how it can be used inline to do bold text, colors, and so on. One way to learn is through right click! This can be used to view the \(\LaTeX\) codes that were used to make things. Some examples of the art and my right clicking: http://youtu.be/KdsRP1G3vpQ

OpenStudy (e.mccormick):

@Miracrown Da V D O stuffs.

Miracrown (miracrown):

Thanks for this @e.mccormick – It has everything that I was after. :^)

OpenStudy (anonymous):